A 500.0 mL cylinder in a car engine was fully filled with gas at 755 mmHg. Starting cold, the piston was pushed to compress the gas to 68.8 mL. The engine did not start and the temperature remained the same, what is the pressure in the cylinder? Answer in 3 sig-dig number without unit mmHg



1
Expert's answer
2021-03-22T06:32:31-0400

"\frac{P_1V_1}{T_1} = \frac{P_2V_2}{T_2}"

Temperature is constant

P1V1 = P2V2


755mmHg × 500 mL = P2×68.8mL


P2 = "\frac{755×500}{68.8}" = 5551.47mmHg


P2 = 555.147×102 mmHg